A person drives back and forth. They start 20 meters away from you at a pace of 10 m/s for 3 seconds, then stand still for 2 seconds, then drive back towards you at a pace of 20 m/s for 3 seconds.
1
Expert's answer
2020-10-27T11:19:41-0400

As here question is incomplete, so i am assuming that we have to find the total time to cover 20 reach to the person.

In the first move, person is moving away, so distance move "=10\times 3= 30m"

Distance covered in the return journey"=20\times3=60 m"

So, in the first move, net distance covered "=(60-30)m=30m"

So, again he will move in forward direction, so time taken to cover 30 m at 20m/sec speed "=\frac{30}{20}s =1.5 sec"

Hence total time taken to reach to the person who is standing"=5+1.5 = 6.5 s"